OpenStudy (anonymous):

I'm in H.S. Calculus & need help. How would you solve for dy/dx using implicit differentiation (with the given points) within the equation: ((X)^1/2)+((y)^1/2)=(2y^2); (1,1)

OpenStudy (anonymous):

\[\huge \frac{1}{2\sqrt{x}}+\frac{y'}{2\sqrt{y}}=4yy'\] is a start

OpenStudy (anonymous):

@Kathy164 will explain why this is the correct first step

OpenStudy (anonymous):

lol damn putting me on the spot like that

OpenStudy (anonymous):

that's ok looks like @juanlopez7 left anyways

OpenStudy (anonymous):

the derivative of \[\sqrt{x}=\frac{ 1 }{ 2\sqrt{x} }\], then you take the derivative of y with respect to x

OpenStudy (anonymous):

so would you subtract \[1/(2\sqrt{y})\] to get the ys alone?

OpenStudy (anonymous):

yea